OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? -b b^2 - 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x^2 + 7x + 3 = 0

OpenStudy (anonymous):

I do not understand what its asking for its a written thing apparently its asking for a new equation?

OpenStudy (anonymous):

b^2 - 4ac If this (called the DISCRIMINANT) equals a positive, perfect square, then the whole standard form equation can be solved using integer factors. b^2 = 7^2 = 49 4ac = 4*2*3 = 24 49-24 = 25 ------> perfect square of 5! so we can bother trying to factor into: (2x+1)(x+3) = 0 and the roots would be x = -1/2 and x = -3